OpenStudy (anonymous):

What is the radius of an automobile tire that turns with a frequency of 11Hz and has a tangential speed of 20 m/s?

OpenStudy (ghazi):

well, tangential velocity= r*w and w= 2*pi *f \[V= r* \omega\] and \[\omega = 2 \pi f\]

OpenStudy (ghazi):

\[\omega = 2*3.14*11=69.08 rad/s\]

OpenStudy (ghazi):

now i hope you can find r

OpenStudy (ghazi):

\[r=\frac{ v }{ \omega }=\frac{ 20 }{ 69.08 }=0.28951939 m =0.29m\]